OpenStudy (anonymous):

2y - 3x = 8 in slope intercept form ?

OpenStudy (amistre64):

what does the slope intercept form look like?

OpenStudy (anonymous):

y = mx + b

OpenStudy (amistre64):

very good; now what do you think should do first to get this to look like that?

OpenStudy (anonymous):

move one of those numbers to become the y ?

OpenStudy (amistre64):

yes... lets "move" the -3x over by adding +3x to each side: 2y = 3x+8 now what can we do?

OpenStudy (anonymous):

isolate the y ?

OpenStudy (amistre64):

yes.. we can do that by "dividing" both sides by 2 right? will that help us?

OpenStudy (anonymous):

yes so . 2y = 3x + 8 y = 3x + 4 ?

OpenStudy (amistre64):

very close.... I see you divided 8 by 2 = 3; but we also need to divide 3x by 2...the number gets us a fraction. y = (3/2)x + 4 do you see that?

OpenStudy (amistre64):

ack!! 8/2 = 4 not 3 lol

OpenStudy (anonymous):

Lol ! so am i getting closer ?

OpenStudy (amistre64):

you are :) does y = (3/2)x + 4 look like: y = mx + b?

OpenStudy (anonymous):

why , yes . yes it does !

OpenStudy (amistre64):

then we are done :)

OpenStudy (anonymous):

thank you ! :D

OpenStudy (anonymous):

Y = 3/2x + 4